OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -5x - y -2 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (anonymous):

5 - Y <= -2

OpenStudy (mertsj):

solid or dotted?

OpenStudy (anonymous):

The inequality sign does NOT have a line under it. So the line isn't solid, so it must be dashed. Meaning its not \[\ge \le \]